OpenStudy (anonymous):

What is the solution to the equation -4(2x+3)=2x+6-(8x+2) ?

OpenStudy (anonymous):

Distribute the -4 on the left side then distribute the negative inside the parentheses on the left side. This should get you -8x-12=2x+6-8x-2. then combine like terms

OpenStudy (anonymous):

A.x = –10 B.x = –8 C.x = -5/2 D.x = -4/7

OpenStudy (anonymous):

my best guess is D?

OpenStudy (anonymous):

is this on connexus?

OpenStudy (anonymous):

Idek what that means but can someone give me the answer as in A,B,D OR C

OpenStudy (anonymous):

????????????????????????????????????????????????????????????????????????????????????????????????????????

OpenStudy (anonymous):

\[-4(2x + 3) = 2x + 6 -(8x + 2)\] \[-8x - 12 = 2x + 6 - 8x - 2\] \[2x = -16\] \[x = -8\] The answer is B.